OpenStudy (anonymous):

Simplify 6q - 3 - 2q. A) 3 - 2q B) 3q - 4 C) 4q - 3 D) q

OpenStudy (anonymous):

\[6q-3-2q=q(6-2)-3=4q-3\]Just combine like terms :-)

OpenStudy (anonymous):

C right?

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

C indeed.

OpenStudy (anonymous):

thanks